I need a real-life application of a linear function. State the application, give the equation of the linear function, and state what the x and y in the application represent. I have to choose at least two values of x to input into my function and find the corresponding y for each. State, in words, what each x and y means in terms of your real-life application.
Answer by elima(1433)

You can rent an electricity generator for $4.80 per hour plus $100.00 for a full tank of petrol.
Let y represent the cost of using the generator.
Let x represent the hours.
The cost is $4.80 per hour plus $100.00 for a the petrol tank.
Translate the statement to an equation.
y = 100.00 + 4.80x
Then a function;
f(x) = 100.00 + 4.80x
Lets choose 6 for the hours,
substitute 6 for x.
f(6) = 100.00 + 4.80 (6)
f(6) = 100.00 + 28.80 = 128.80
So, the cost of renting the generator for 6 hours is $128.80
Now choose another value for x and solve equation just as I did for y.
:)